Suppose M = 8, P1 = 4, P2 = 2. Graph the budget set for this consumer.

b) Define the opportunity cost of good 1 in terms of good 2.

c) Determine the opportunity cost of good 1 at the point where 21 = 1.

d) Does the opportunity cost of 1 change as we move along the budget line? Explain.

e) Now suppose that M = 16, P1 = 2, P2 = 4. Graph the consumer's budget set in this case

f) Can we rank these budget sets in terms of which is better for the consumer? Explain.


1
Expert's answer
2021-09-23T09:12:45-0400

a)

Given m=8,p1=4 and P2=2

Budget will be "8=4x_1+2x_2"



If x1=0 "X_2=\frac{8}{2}=4"

If x2=0 "X_1=\frac{8}{4}=2"


b)

Opportunity cost of good 1

"=\frac{\Delta consumption \space of\space good \space 2}{\Delta consumption \space of\space good \space 1}=\frac{4}{2}=2"


c)

Opportunity cost of good 1 when x1=0

"\frac{4-2}{1}=2"


d)

Opportunity cost does not change as we move along the budget line as the budget line is a straight line.


e)

Given m=16,p1=2 and P2=4

Budget will be "16=2x_1+4x_2"



If x1=0 "X_2=\frac{16}{4}=4"

If x2=0 "X_1=\frac{16}{2}=8"

f)

In budget 16=2x1+4x2 consumer can consume more quantity of good 1 and 2 compared to budget 8=4x1+2x2